the perimeter of a rectangle is 60 meters and the area is 221 square meters. Find its dimensions
Answer by lwsshak3(11628)

the perimeter of a rectangle is 60 meters and the area is 221 square meters. Find its dimensions
***
let x=length of rectangle
let y=width of rectangle
2x+2y=perimeter
2x+2y=60
x+y=30
y=30-x
length*width=area
xy=221
x(30-x)=221
30x-x^2=221
x^2-30x+221=0
(x-13)(x-17)
x=13
or
x=17
dimensions:
length=17 m
width=13 m